# Leveraging Large Language Models for Collective Decision-Making

### v2

@@ 6.3. Effect of Incorporating the Knowledge Graph
This pattern is as expected since introducing multiple additional constraints embedded in the complexity of the knowledge graph creates more challenges in balancing the members’ preferences. For example, suppose that one of the members must attend a meeting. Any options must accommodate this requirement, which may, in turn, yield options that are not satisfactory or require compromises (i.e., still feasible but with fewer preferences satisfied) for the other members.
@@ 6.4. Comparison with Baselines and Robustness Checks
Finally, in Table 2 we found that the results remain almost unchanged after repeatedly paraphrasing the schedule preferences. This suggested our system was robust to paraphrased preferences.
@@ 7.2. Measures and Analysis
In the survey questions, we used a 5-point Likert scale ranging from 1 to 5, with 1 representing the poorest performance and 5 representing the best rating. We also used free-text questions and asked the participants for their open-ended feedback on their evaluation.
For each question that elicits numerical ratings, we computed the average of the participants’ responses in two ways: We calculated the average over all of the participants and the average for each group. For each of the averages, we calculated a 95% confidence interval (two-sided) in our system using the Student-t distribution with 44 degrees of freedom for all participants, 13 degrees of freedom for Group A, 12 degrees of freedom for Group B, and 17 degrees of freedom for Group C.
@@
Question
Average Rating
@@
To what degree do you believe that the system takes your preferences into account when presenting options?
3.51 $\pm$ 0.35
Do you find the reasons behind the options acceptable based on your preferences?
3.32 $\pm$ 0.37
After seeing all members’ individual preferences, to what degree do you believe that our system takes your preferences into account when presenting options?
3.54 $\pm$ 0.35
After seeing all members’ individual preferences, do you find the reasons behind the options acceptable?
3.46 $\pm$ 0.36
How well do you agree the options presented reflect the preferences of the members?
3.07 $\pm$ 0.32
How confident are you in using our system to coordinate members’ preferences in meeting scheduling?
3.49 $\pm$ 0.33
@@
To what degree do you believe that the system takes your preferences into account when presenting options?
3.64 $\pm$ 0.59
Do you find the reasons behind the options acceptable based on your preferences?
3.36 $\pm$ 0.59
After seeing all members’ individual preferences, to what degree do you believe that our system takes your preferences into account when presenting options?
3.91 $\pm$ 0.60
After seeing all members’ individual preferences, do you find the reasons behind the options acceptable?
4.09 $\pm$ 0.53
How well do you agree the options presented reflect the preferences of the members?
3.36 $\pm$ 0.66
How confident are you in using our system to coordinate members’ preferences in meeting scheduling?
4.09 $\pm$ 0.53
@@
To what degree do you believe that the system takes your preferences into account when presenting options?
3.17 $\pm$ 0.73
Do you find the reasons behind the options acceptable based on your preferences?
3.08 $\pm$ 0.80
After seeing all members’ individual preferences, to what degree do you believe that our system takes your preferences into account when presenting options?
3.67 $\pm$ 0.70
After seeing all members’ individual preferences, do you find the reasons behind the options acceptable?
3.50 $\pm$ 0.76
How well do you agree the options presented reflect the preferences of the members?
3.08 $\pm$ 0.61
How confident are you in using our system to coordinate members’ preferences in meeting scheduling?
3.33 $\pm$ 0.75
@@
To what degree do you believe that the system takes your preferences into account when presenting options?
3.67 $\pm$ 0.60
Do you find the reasons behind the options acceptable based on your preferences?
3.44 $\pm$ 0.65
After seeing all members’ individual preferences, to what degree do you believe that our system takes your preferences into account when presenting options?
3.22 $\pm$ 0.59
After seeing all members’ individual preferences, do you find the reasons behind the options acceptable?
3.06 $\pm$ 0.59
How well do you agree the options presented reflect the preferences of the members?
2.89 $\pm$ 0.55
How confident are you in using our system to coordinate members’ preferences in meeting scheduling?
3.22 $\pm$ 0.51
@@ 8.1. Interpreting the Results
Our evaluation results demonstrated that our system excels across a wide array of metrics. Specifically, the system has shown positive performance in addressing all the posed research questions.
@@
Second, in terms of effectiveness (RQ2), most of the experiments showed consistent progression in the satisfaction ratio of the decision candidate (cf. Figures 6, 4 and 5). This outcome demonstrated that the coordination module was successful in its task of progressively refining its proposed options to satisfy more members. The effectiveness of the system was also exemplified in scenarios where a meeting involved many members and finding a mutually agreeable meeting time was challenging (cf. $n=5$ in Figure 5). We noted that adopting a smaller number of options $K$ achieved higher satisfaction ratios, as evidenced in Figure 5. Lastly, when attendance constraints and the knowledge graph were integrated (cf. Figure 6), the system continued to improve the satisfaction ratio, albeit at a significantly reduced magnitude. This suggested that despite the added heterogeneities and complexity, the system was still able to accommodate members’ preferences, though its performance was less effective.
Third, for the quality of proposed options (RQ3), our findings showed that the satisfaction scores of the decision candidate were improved across the settings. In particular, we observed that as the number of members increased (cf. Figure 4), there was a large improvement in the percentage change in the satisfaction score. On the other hand, when the number of options was varied from $K=2$ to $K=3$ or $4$, the satisfaction scores were reduced (cf. Figure 5). It was inconclusive whether the options that were successively proposed by the coordination had lower equity scores.
@@ 8.2. Ablation Study
Firstly, as shown in Table 1, the intent extraction module is essential to the functionality of the system both in terms of user-friendliness as well as system performance. Specifically, we observed that even though with the intent extraction module, the performance was worse in the first round (cf. the performance comparison between the two baselines in Table 1), the effect of this module became positive when the coordination proceeds in rounds. This is because the intent extraction module allows the users to communicate their preferences, express their opinions about the suggested options, or propose alterations to the existing options. Moreover, we found that the system was robust to paraphrased preferences, indicating that the intent extraction module effectively extracted the same intents from similarly phrased preferences.
@@
Thirdly, the positive assessment of the system by human participants aligned with the scores assigned by the evaluation module, showcasing the effectiveness of the evaluation module in measuring satisfaction.
Finally, including the knowledge graph could capture more complexities and heterogeneities (e.g., social dynamics, responsibilities, etc.) for the decision-making process; however, this also had negative impacts on the performance metrics, as discussed in Section 6.
@@ 8.3. Challenges and Limitations in System Implementation
The first set of challenges involves the token limits, the high number of requests the system sends to the LLMs, and the latency of processing the language input when the meeting size is large. A potential solution could be calibrating the coordination protocol. In the current design, the coordination module must elicit opinions from all of the members at every round. A potential enhancement could entail the coordination module selecting a subset of the members for necessary conversation at each round, guided by the previous conversation history and proposed options. This approach carries direct benefits on two sides: It reduces the frequency of calls to the LLM and shortens the size of the prompt request (as it includes the conversation with fewer members). These reductions would further translate to smaller processing latency and system load for the LLMs. When the system is deployed on a large scale, the savings on operational costs could be substantial.
@@
Specifically, in the context of scheduling meetings, the availability of each member’s work calendar is another practical factor taken into consideration. In fact, it is not straightforward to expand our system to incorporate this factor into the LLM’s reasoning process. To examine the LLM’s ability to incorporate work calendars into its reasoning process, we conducted an experiment in which constraints based on each member’s availability were mathematically encoded into the LLM prompt for the respective member. We observed that the LLM’s performance degraded, suggesting that LLMs are less capable of dealing with hard constraints encoded in mathematical format (see, e.g., (Wei et al., 2022; Rae et al., 2021) and the references therein). Two potential solutions for this challenge are using an external tool that is tailored to understand the availability of the members or fine-tuning the model to handle the members’ availability constraints.
@@ 8.4. Privacy Concerns
A noteworthy concern related to our system is the management of information leakage from members. Specifically, it’s crucial to prevent the exposure of sensitive information to other members, such as undisclosed preferences, sensitive attributes, and confidential details from the members’ past behavior (which can be, e.g., exposed via the knowledge graph). Therefore, enhancing the coordination module’s capacity to protect sensitive information becomes essential to the success of the system. Such privacy concerns have gained prominence (cf. (Kandpal et al., 2022; Carlini, 2020; Brown et al., 2022) and the references therein), and various strategies have been proposed to address the concern. Such mitigation strategies can include implementing LLM-based redaction and filtering modules and fine-tuning the LLMs to preserve privacy. In the context of our model, it is certainly possible to instruct the intent extraction prompts to redact any potentially sensitive information or to introduce noise in the preferred preferences (i.e., slightly perturb the reported preferences) before sending them to the coordination module.
@@ 8.5. A New Way of Work Coordination
Our system exhibits the potential to facilitate collective decision-making by coordinating conversations between members. It actively understands members’ preferences and suggests alternatives that aid in reaching consensus across various contexts, such as meeting scheduling, work plan formulation, idea brainstorming, and discussion reconciliation. The use of LLMs can enhance productivity and creativity by expediting the decision-making process through real-time analysis of vast data and presentation of relevant insights. Furthermore, LLMs consider multiple perspectives, ensuring all relevant viewpoints are taken into account for more inclusive and comprehensive decision-making. Beyond facilitating collective decision-making, LLMs can significantly boost productivity by alleviating non-essential work, thereby saving time and effort and reducing costs.
@@ 9.1. Collective Decision-Making in Work Environments
The growing trend of organizations becoming more distributed, coupled with the necessity for collaboration among employees in various locations, has led to a wide range of open-ended preferences and organizational asymmetries (Mok et al., 2023; Reinecke et al., 2013). The trend prompts the development of distributed meeting scheduling assistants, such as Calendar.help (Cranshaw et al., 2017). However, as mentioned in Section 2, Calendar.help considers workers’ calendars and cannot handle open-ended preferences, whereas our system can identify meeting schedules that balance diverse open-ended preferences among several members considering social relations and organizational asymmetries.
@@ 9.2. Understanding Human Preferences Using Large Language Models
Large language models have been utilized in many studies to comprehend human preferences. These models analyze historical interactions and human feedback to discern preferences for specific topics, products, or services. A notable application of LLMs is their use in tasks such as collaborative filtering to understand human preferences. The research by (Kang et al., 2023) explores whether LLMs can learn to rank similarly to ranking algorithms in a zero-shot or few-shot manner and shows that LLMs can achieve performance akin to collaborative filtering algorithms with substantially less data. Similarly, in (Sanner et al., 2023) it is shown that zero-shot and few-shot-trained LLMs are able to perform competitively compared to state-of-the-art cold-start collaborative filtering methods for recommendations from language and item-based preferences, offering additional explainability over traditional methods. Our work contributes to this field of work as our system extracts group members’ individual preferences from their dialogues with the system and accordingly makes suggestions for the group.
Another development is using machine learning methods to generate consensus statements by reconciling various opinions. The work of (Cao et al., 2018) introduces AGREE, which is a neural collaborative filtering method for groups and users by utilizing a learned preference aggregation mechanism based on attention. Compared to AGREE, our method is also able to aggregate different preferences; however, in our case, we do not specifically train models for handling preferences, and, instead, we utilize LLMs to consolidate them. Besides, in (Bakker et al., 2022) the authors fine-tune LLMs to aggregate members’ opinions and generate consensus statements. They train reward models to predict individual preferences, enabling them to quantify and rank consensus statements according to different social choice functions. Our work contributes to this development as it also leverages LLMs to find consensus, balancing the opinions of multiple members.
@@ 9.3. Simulating Human Behavior Using Large Language Models
Particularly, simulating agents with human-like behavior using LLMs has garnered significant interests in the field of social sciences (Grossmann et al., 2023). For instance, LLM-simulated agents have been employed to study social behavior (Park et al., 2023), replicate human subject studies (Aher et al., 2022), perform market research (Brand et al., 2023), act as economic agents (Horton, 2023), conduct qualitative interviews at a large scale (Chopra and Haaland, 2023), and perform crowdsourcing tasks (Veselovsky et al., 2023). Our paper contributes to this line of study by utilizing LLM-simulated members to test and evaluate our system, addressing the challenges and costs associated with conducting user studies involving human participants (e.g., the length of 18 months for the user study of Calendar.help (Cranshaw et al., 2017)).
@@ 9.5. AI and Collective Decision-Making
Recent research has explored how artificial intelligence can assist in collaborative decision-making, with roots tracing back to graph mining techniques, see, e.g., (Zhang et al., 2018; Gu et al., 2021). The work of (Bakker et al., 2022) demonstrates that fine-tuning LLMs can generate statements that maximize the expected approval from a diverse group of people with diverse opinions.Our research contributes to the growing area of using machine learning methods for decision making (cf. (Landemore, 2020; Fishkin, 2009)).
@@ 11. Ethics Statement
Due to our evaluation involving participants assessing the system’s ability to exhibit human-like reasoning (RQ4), we are cognizant of potential privacy concerns for the individuals in our study. To address these concerns, our research underwent a formal privacy review within our organization, ensuring that individual employees’ identities remain protected. Additionally, all survey participants provided explicit consent for us to use their responses in our research, including this paper. You can find the consent form used for our study in Appendix D. We have taken measures to anonymize all results by removing any names or email addresses and presenting responses at a higher-level aggregation. Individual data has not been analyzed or disclosed in this study. Furthermore, we have omitted potentially sensitive metadata, such as indicators of seniority and job function. Access to the dataset has been limited to the research team involved in this project to maintain confidentiality. Participants were compensated for their participation in the user study.

### v3

@@ 6.3. Effect of Incorporating the database
This pattern is expected since introducing multiple additional constraints embedded in the database’s complexity creates more challenges in balancing the members’ preferences. For example, suppose that one of the members must attend a meeting. Any options must accommodate this requirement, which may, in turn, yield options that are not satisfactory or require compromises (i.e., still feasible but with fewer preferences satisfied) for the other members.
@@ 6.4. Comparison with Baselines and Robustness Checks
Finally, in Table 2 we found that the results remain almost unchanged after repeatedly paraphrasing the schedule preferences. This suggested our system was robust to the paraphrased preferences.
@@ 7.2. Measures and Analysis
In the survey questions, we used a 5-point Likert scale ranging from 1 to 5, with 1 representing the poorest performance and 5 representing the best rating. We also used free-text questions and asked the participants for their free-text feedback on their evaluation.
For each question that elicits numerical ratings, we computed the average of the responses of the participants in two ways: We calculated the average for all participants and the average for each group. For each of the averages, we calculated a 95% confidence interval (two-sided) in our system using the Student-t distribution with 44 degrees of freedom for all participants, 13 degrees of freedom for Group A, 12 degrees of freedom for Group B, and 17 degrees of freedom for Group C.
@@
Question Average Rating
@@
To what degree do you believe that the system takes your preferences into account when presenting options? 3.51 $\pm$ 0.35
Do you find the reasons behind the options acceptable based on your preferences? 3.32 $\pm$ 0.37
After seeing all members’ individual preferences, to what degree do you believe that our system takes your preferences into account when presenting options? 3.54 $\pm$ 0.35
After seeing all members’ individual preferences, do you find the reasons behind the options acceptable? 3.46 $\pm$ 0.36
How well do you agree the options presented reflect the preferences of the members? 3.07 $\pm$ 0.32
How confident are you in using our system to coordinate members’ preferences in meeting scheduling? 3.49 $\pm$ 0.33
@@
To what degree do you believe that the system takes your preferences into account when presenting options? 3.64 $\pm$ 0.59
Do you find the reasons behind the options acceptable based on your preferences? 3.36 $\pm$ 0.59
After seeing all members’ individual preferences, to what degree do you believe that our system takes your preferences into account when presenting options? 3.91 $\pm$ 0.60
After seeing all members’ individual preferences, do you find the reasons behind the options acceptable? 4.09 $\pm$ 0.53
How well do you agree the options presented reflect the preferences of the members? 3.36 $\pm$ 0.66
How confident are you in using our system to coordinate members’ preferences in meeting scheduling? 4.09 $\pm$ 0.53
@@
To what degree do you believe that the system takes your preferences into account when presenting options? 3.17 $\pm$ 0.73
Do you find the reasons behind the options acceptable based on your preferences? 3.08 $\pm$ 0.80
After seeing all members’ individual preferences, to what degree do you believe that our system takes your preferences into account when presenting options? 3.67 $\pm$ 0.70
After seeing all members’ individual preferences, do you find the reasons behind the options acceptable? 3.50 $\pm$ 0.76
How well do you agree the options presented reflect the preferences of the members? 3.08 $\pm$ 0.61
How confident are you in using our system to coordinate members’ preferences in meeting scheduling? 3.33 $\pm$ 0.75
@@
To what degree do you believe that the system takes your preferences into account when presenting options? 3.67 $\pm$ 0.60
Do you find the reasons behind the options acceptable based on your preferences? 3.44 $\pm$ 0.65
After seeing all members’ individual preferences, to what degree do you believe that our system takes your preferences into account when presenting options? 3.22 $\pm$ 0.59
After seeing all members’ individual preferences, do you find the reasons behind the options acceptable? 3.06 $\pm$ 0.59
How well do you agree the options presented reflect the preferences of the members? 2.89 $\pm$ 0.55
How confident are you in using our system to coordinate members’ preferences in meeting scheduling? 3.22 $\pm$ 0.51
@@ 8.1. Interpreting the Results
Our evaluation results demonstrated that our system excels across a wide range of metrics. Specifically, it has performed well in addressing all the research questions posed.
@@
Second, in terms of effectiveness (RQ2), most of the experiments showed consistent progression in the satisfaction ratio of the decision candidate (cf. Figures 6, 4 and 5). This outcome demonstrated that the coordination module was successful in its task of progressively refining its proposed options to satisfy more members. The effectiveness of the system was also exemplified in scenarios where a meeting involved many members and finding a mutually agreeable meeting time was challenging (cf. $n=5$ in Figure 5). We noted that adopting a smaller number of options $K$ achieved higher satisfaction ratios, as evidenced in Figure 5. Lastly, when attendance constraints and the database were integrated (cf. Figure 6), the system continued to improve the satisfaction ratio, albeit at a significantly reduced magnitude. This suggested that despite the added heterogeneities and complexity, the system was still able to accommodate members’ preferences, though its performance was less effective.
Third, regarding the quality of proposed options (RQ3), our findings showed that the decision candidate’s satisfaction scores improved across the settings. In particular, we observed that as the number of members increased (cf. Figure 4), there was a large improvement in the percentage change in the satisfaction score. On the other hand, when the number of options was varied from $K=2$ to $K=3$ or $4$, the satisfaction scores were reduced (cf. Figure 5). It was inconclusive whether the options that were successively proposed by the coordination had lower equity scores.
@@ 8.2. Ablation Study
[…]
@@
Thirdly, human participants’ positive assessments of the system aligned with the scores assigned by the evaluation module, showcasing the effectiveness of the evaluation module in measuring satisfaction.
Finally, including the database could capture more complexities and heterogeneities (e.g., social dynamics, responsibilities, etc.) for the decision-making process; however, this also had negative impacts on the performance metrics, as discussed in Section 6.
@@ 8.3. Challenges and Limitations in System Implementation
[…]
@@
[…]
We observed a decline in performance in our experiment when incorporating the database that includes the social graph and employees’ backgrounds. While LLMs can handle structural information to some extent, they do not excel at it. Two approaches could be applied to address this challenge: augmenting LLMs with external functions that assist in managing network properties or fine-tuning LLMs for this task using more targeted data.
Specifically, when scheduling meetings, the availability of each member’s work calendar is another practical factor to consider. In fact, expanding our system to incorporate this factor into the LLM’s reasoning process is not straightforward. To examine the LLM’s ability to incorporate work calendars into its reasoning process, we conducted an experiment in which constraints based on each member’s availability were mathematically encoded into the LLM prompt for the respective member. We observed that the LLM’s performance degraded, suggesting that LLMs are less capable of dealing with hard constraints encoded in mathematical format (see, e.g., (Wei et al., 2022; Rae et al., 2021) and the references therein). The two approaches mentioned—augmenting LLMs and fine-tuning—can also be applied here to enhance the system’s ability to process calendar data. For instance, in (Lawless et al., 2024), LLMs are used to translate free-text constraints into code, which is then processed by an optimization package to identify suitable meeting times.
@@ 8.4. Privacy Concerns
[…]
Our system uses private communication between each member and the system, enabling members to freely express their preferences regarding group decisions. However, there remains a concern that private information about a member—such as undisclosed preferences, sensitive attributes, or confidential details from past behaviors (potentially accessible via the database)—could be inadvertently revealed to the group when the system provides explanations for its suggestions. Therefore, enhancing the coordination module’s capacity to protect sensitive information becomes essential to the success of the system. Such privacy concerns have gained prominence (cf. (Kandpal et al., 2022; Carlini, 2020; Brown et al., 2022) and the references therein), and various strategies have been proposed to address the concern. Such mitigation strategies can include implementing LLM-based redaction and filtering modules and fine-tuning the LLMs to preserve privacy. In the context of our model, it is certainly possible to instruct the intent extraction prompts to redact any potentially sensitive information or to introduce noise in the preferred preferences (i.e., slightly perturb the reported preferences) before sending them to the coordination module.
[…]
8.5. LLM-simulated Users versus Human Users
We acknowledge that human behavior cannot be fully captured by LLM-simulated users. We list several key differences as follows: As mentioned before, humans may exhibit behavioral inconsistencies, such as changing preferences or hesitating to make definitive choices—traits that are difficult to simulate accurately with LLMs. Real users bring emotional and psychological factors into group decision-making, including motivation, fatigue, social influence, and stress under time constraints. Additionally, human users may have privacy and trust concerns, particularly if they worry that sensitive information could be exposed or misinterpreted by the system. This may affect their willingness to share authentic preferences, a factor typically absent in LLM simulations. Finally, human users may employ strategies to influence group decisions to their advantage, an aspect commonly observed in social interactions. For example, they may misrepresent their preferences to prompt the coordinator to suggest a more favorable option. LLMs struggle to capture such complex, strategic thinking.
@@ 8.6. A New Way of Work Coordination
[…]
@@ 9.1. Collective Decision-Making in Work Environments
The growing trend of organizations becoming more distributed, coupled with the necessity for collaboration among employees in various locations, has led to a wide range of free-text preferences and organizational asymmetries (Mok et al., 2023; Reinecke et al., 2013). The trend prompts the development of distributed meeting scheduling assistants, such as Calendar.help (Cranshaw et al., 2017). However, as mentioned in Section 2, Calendar.help considers workers’ calendars and cannot handle free-text preferences, whereas our system can identify meeting schedules that balance diverse free-text preferences among several members, considering social relations and organizational asymmetries.
@@ 9.2. Understanding Human Preferences Using Large Language Models
Large language models have been utilized in many studies to comprehend human preferences. These models analyze historical interactions and human feedback to discern preferences for specific topics, products, or services. A notable application of LLMs is their use in tasks such as collaborative filtering to understand human preferences. The research by (Kang et al., 2023) explores whether LLMs can learn to rank similarly to ranking algorithms in a zero-shot or few-shot manner and shows that LLMs can achieve performance akin to collaborative filtering algorithms with substantially less data. Similarly, in (Sanner et al., 2023), it is shown that zero-shot and few-shot-trained LLMs are able to perform competitively compared to state-of-the-art cold-start collaborative filtering methods for recommendations from language and item-based preferences, offering additional explainability over traditional methods. Our work contributes to this field of work as our system extracts group members’ individual preferences from their dialogues with the system and accordingly makes suggestions for the group.
Another development is using machine learning methods to generate consensus statements by reconciling various opinions. The work of (Cao et al., 2018) introduces AGREE, which is a neural collaborative filtering method for groups and users by utilizing a learned preference aggregation mechanism based on attention. Compared to AGREE, our method is also able to aggregate different preferences; however, in our case, we do not specifically train models for handling preferences, and instead, we utilize LLMs to consolidate them. Besides, in (Bakker et al., 2022), the authors fine-tune LLMs to aggregate members’ opinions and generate consensus statements. They train reward models to predict individual preferences, enabling them to quantify and rank consensus statements according to different social choice functions. Our work contributes to this development as it also leverages LLMs to find consensus, balancing the opinions of multiple members.
@@ 9.3. Simulating Human Behavior Using Large Language Models
Particularly, simulating agents with human-like behavior using LLMs has garnered significant interest in the field of social sciences (Grossmann et al., 2023). For instance, LLM-simulated agents have been employed to study social behavior (Park et al., 2023), replicate human subject studies (Aher et al., 2022), perform market research (Brand et al., 2023), act as economic agents (Horton, 2023), conduct qualitative interviews at a large scale (Chopra and Haaland, 2023), and perform crowdsourcing tasks (Veselovsky et al., 2023). Our paper contributes to this line of study by utilizing LLM-simulated members to test and evaluate our system, addressing the challenges and costs associated with conducting user studies involving human participants (e.g., the length of 18 months for the user study of Calendar.help (Cranshaw et al., 2017)).
@@ 9.5. AI and Collective Decision-Making
Recent research has explored how artificial intelligence can assist in collaborative decision-making, with roots tracing back to graph mining techniques, see, e.g., (Zhang et al., 2018; Gu et al., 2021). The work of (Bakker et al., 2022) demonstrates that fine-tuning LLMs can generate statements that maximize the expected approval from a diverse group of people with diverse opinions. Our research contributes to the growing area of using machine learning methods for decision-making (cf. (Landemore, 2020; Fishkin, 2009)).
@@ 11. Ethics Statement
Due to our evaluation involving participants assessing the system’s ability to exhibit human-like reasoning (RQ4), we are cognizant of potential privacy concerns for the individuals in our study. To address these concerns, our research underwent a formal privacy review within our organization, ensuring that individual employees’ identities remain protected. Additionally, all survey participants provided explicit consent for us to use their responses in our research, including this paper. You can find the consent form used for our study in Appendix D. We have taken measures to anonymize all results by removing any names or email addresses and presenting responses at a higher-level aggregation. Individual data has not been analyzed or disclosed in this study.
Furthermore, we have omitted potentially sensitive metadata, such as indicators of seniority and job function. To maintain confidentiality, access to the dataset has been limited to the research team involved in this project. Participants were compensated for their participation in the user study.